Mr. Wallace can paint a fence in 2 hours. His son Greg can paint the fence in 6 hours. Mr. Wallace painted alone for 1 hour and stopped working. How many hours did it take Greg to finish the job? Show your work.
Answer by CharlesG2(834)

Mr. Wallace can paint a fence in 2 hours. His son Greg can paint the fence in 6 hours. Mr. Wallace painted alone for 1 hour and stopped working. How many hours did it take Greg to finish the job? Show your work.


Mr. Wallace can paint fence in 2 hours and worked alone for 1 hour and stopped,
so that means he did half the job, would take if he worked alone 1 more hour to finish the job.
Greg can paint the whole fence in 6 hours but only needs to paint half of the fence, working alone that should take Greg 3 hours.

The problem described seems to imply Greg did the remainder of the fence alone. It also said Mr. Wallace worked alone for 1 hour and could have done the whole fence in 2 hours, so Mr. Wallace did half the fence, leaving half the fence left.

I might be in error, if this problem involved working together for the 2nd half of the fence then:
T time = 1/(1/W + 1/G), W = Mr. Wallace time in hours = 1,
G = Greg time in hours = 3
the units are: hrs = job/(job/hrs + job/hrs) = job * hrs/job = hrs
T = 1/(1 + 1/3) = 1/(4/3) = 3/4 hrs = 45 minutes if worked together
(60 minutes in an hour, 3/4 of 60 is 45)